OpenStudy (anonymous):

Help please!! How do I find the derivative of y=e^(sin2theta)?

OpenStudy (ipwnbunnies):

Is the power (sin(theta))^2 or sin(2*theta)

OpenStudy (anonymous):



OpenStudy (ipwnbunnies):

Cool, do you know the general form of the derivative of e? \[\frac{d}{dx}[e^u] = e^u * \frac{du}{dx}\]

OpenStudy (ipwnbunnies):

You rewrite the function, then multiply it by the derivative of the power.

OpenStudy (anonymous):

Yeah, basically the derivative of e to whatever power is just e, right?

OpenStudy (ipwnbunnies):

What? o.o That's not what I said.

OpenStudy (ipwnbunnies):

Like, what's the derivative of e^x?

OpenStudy (anonymous):

I'm a little confused...

OpenStudy (ipwnbunnies):

Do you know what the derivative of e^x is at least?

OpenStudy (anonymous):

It's e^x, right?

OpenStudy (ipwnbunnies):

Right, good. So what would e^(2x) be?

OpenStudy (anonymous):

e^2x

OpenStudy (anonymous):

?

OpenStudy (ipwnbunnies):

No, the derivative will be e^(2x) times the derivative of (2x).

OpenStudy (anonymous):

OH, because of the chain rule?